Q / Construct the difference table from the following data: * 1 3 5 7 9 11 and then find f(6.2)? y 10 4 5 20 -7 13
Answer #1

Newtoris Forward difference table y 생 By 행 sly 5g l0 - -6 3 니 구 서 (디 263 5 5 56 208 115 -42 20 7 27 89 니가 9 20 13 Newtons foYG) = -1.528 + 15.23x² - 66.018² + 117.228–54.98 Now, substitute x = 6-2 4(6-2) = -1.52% 6.24 +15:23 X6-33-66.01x62? +117.27